OpenStudy (anonymous):

divide (X+5)(X-3)/3(X+5) (by X+5)

OpenStudy (anonymous):

You have a common X+5 term that can cancel out

OpenStudy (anonymous):

isit x-3/3

OpenStudy (anonymous):

I am a bit confused is your fraction this \[\frac{(x+5)(x-3)}{3(x+5)}\] in the beginning?

OpenStudy (anonymous):

yes, and they ask to divide by x+5

OpenStudy (anonymous):

ok so the (X+5) term will drop out and then you will have \[\frac{\frac{(x-3)}{3}}{(X+5)}\] you can multiply by the reciprocal when you have a fraction within a fraction \[\frac{(X-3)}{3}*\frac{(X+5)}{1}\]

OpenStudy (anonymous):

-2x+5/3

OpenStudy (anonymous):

You have to use FOIL to multiply the top terms (X-3)(X+5) ... the bottom term is correct

OpenStudy (anonymous):

oh x^2+2x-15/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can simply a little more and distribute the bottom 3 into each of the top terms numbers

OpenStudy (anonymous):

x^2+2x-5

OpenStudy (anonymous):

the whole top term is divided by 3 not just the -15

OpenStudy (mertsj):

\[\frac{(x+5)(x-3)}{3(x+5)}\times \frac{1}{x+5}=\frac{x-3}{3(x+5)}\]

OpenStudy (anonymous):

o yes my apologizes Mertsj is correct

OpenStudy (mertsj):

no

OpenStudy (anonymous):

x-3/3x+15

OpenStudy (mertsj):

\[\frac{x-3}{3x+15}\]

OpenStudy (mertsj):

NO!!!!!

OpenStudy (mertsj):

Only factors may be cancelled.